OpenStudy (anonymous):

If anyone can help me with the steps to factoring trinomials I would really appreciate it!

OpenStudy (bahrom7893):

what is it?

OpenStudy (bahrom7893):

2x^2+11x+5?

OpenStudy (anonymous):

Yeah thats it! I need the steps not just the answer!

OpenStudy (bahrom7893):

D = 11^2 - 4*2*5 = 121 - 40 = 81 x1 = [-11 - sqrt(81)]/(2*2) = [-11-9]/4 = -20/4 = -5 x2 = [-11 + sqrt(81)]/(2*2) = [-11+9]/4 = -2/4 = -1/2

OpenStudy (bahrom7893):

so to factor: (x-x1)(x-x2) = (x+5)(x+1/2)

OpenStudy (bahrom7893):

sorry I don't know how to factor but you can use quadratics and you know how in some cases you can't factor, quadratics will always get u the answer..

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

(2x+1)(x+5)

OpenStudy (anonymous):

you have to think what will add to 11 and multiply to 5

OpenStudy (anonymous):

and you know the first term has to be 2x and the second x

OpenStudy (anonymous):

does that help i can explain more